evaluate ∫ tan^-1(sin 2x / 1 + cos 2x) dx

evaluate
∫  tan^-1(sin 2x / 1 + cos 2x) dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int tan^{-1}(\frac{sin2x}{1+cos2x})dx
I = \int tan^{-1}(\frac{2sinx.cosx}{2cos^{2}x})dx
I = \int tan^{-1}(\frac{sinx}{cosx})dx
I = \int tan^{-1}(tanx)dx
I = \int xdx
I = \frac{x^2}{2} + c